OpenStudy (anonymous):

solve the following inequality 9x-80>9/x

OpenStudy (anonymous):

x>9

OpenStudy (anonymous):

oh hold the phone missing a piece

OpenStudy (anonymous):

need to solve \[9x-80>\frac{9}{x}\] \[9x-80-\frac{9}{x}>0\] \[\frac{(x-9)(9x+1)}{x}>0\] critical points are at -1/9, 0, 9

OpenStudy (anonymous):

solution is \[-\frac{1}{9}<x<0, x>9\] two intevals

OpenStudy (anonymous):

@Luis Rivera you cannot multiply by x, because you do not know whether x is positive or negative

OpenStudy (anonymous):

so you don't know whether to keep the inequality or switch it. you have to put everything on one side of the inequality and make it > 0